# RetroGFN: Diverse and Feasible Retrosynthesis using GFlowNets

### v1

@@ 6 Conclusions
In this paper, we provided empirical arguments for the importance of reporting the round-trip accuracy in the single-step retrosynthesis model evaluation. Leveraging the GFlowNet framework which is designed for tasks where plenty of sensible solutions are desired, we developed a RetroGFN model that achieves competitive results on top-k accuracy and performs outstandingly on the top-k round-trip accuracy. We discuss the limitations of the paper in the Appendix D.
@@ Acknowledgements
The research of P. Gaiński and M. Śmieja was supported by the National Science Centre (Poland), grant no. 2022/45/B/ST6/01117. The research of J. Tabor was supported by the Foundation for Polish Science co-financed by the European Union under the European Regional Development Fund in the POIR.04.04.00-00-14DE/18-00 project carried out within the Team-Net program. The research of M. Koziarski was supported by funding from CQDM Fonds d’Accélération des Collaborations en Santé (FACS) / Acuité Québec and Genentech. We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid (ACK Cyfronet Athena, HPC) for providing computer facilities and support within computational grant no PLG/2023/016550. For the purpose of Open Access, the author has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
@@ Appendix A RetroGFN Details
All neural networks in RetroGFN used the same hidden dimension $h=200$. To obtain initial node and edge features for products, we used featurization from [58] implemented in the DGL library [59]. For the reactant pattern, we used the same edge featurization and a custom node featurization that accounted for atom type, degree, aromaticity, whether the atom was mapped in the original template, relative charge difference between product and reactant atom in the original template, and analogous implicit hydrogen difference. The node features for both products and reactant patterns were enriched with random walk positional encoding [60] of size $\mathrm{n\_random\_walk}=16$.
Product node encoder $\mathrm{gnn}_{1}$ consists of $\mathrm{num\_layer\_1}=4$ layers of the MPNN convolution [61] and one Transformer layer with $\mathrm{num\_heads}=8$. The reactant pattern encoder differs only in the number of layers $\mathrm{num\_layer\_2}=3$. Multi-layer perceptrons $\mathrm{MLP}_{1},\mathrm{MLP}_{2},\mathrm{MLP}_{3}$ had one hidden layer (with hidden dimension $h$) and used the GeLU activation function.
During training, we used a combination of three sampling methods: 1) standard exploratory sampling from the forward policy $P_{F}$ with some $\epsilon$ probability of taking random actions, 2) backward sampling from replay buffer [62, 63, 64], and 3) backward sampling from the dataset $D$. Backward sampling starts with a terminal state and samples the trajectory in the backward direction using the backward policy.
During the training probability of taking random action in the forward policy was set to $\epsilon=0.05$, the number of sampled forward trajectories in the batch was $\mathrm{n\_forward}=16$, and the analogous numbers for backward dataset trajectories and backward replay buffer trajectories were $\mathrm{n\_dataset}=96$ and $\mathrm{n\_replay}=16$. The model was trained with Adam optimizer [65] with a learning rate $lr=0.0005$ (with other parameters set to default values in the torch implementation) for $\mathrm{n_{i}terations}=25000$ iterations. In the evaluation, the forward policy temperature was set to $\alpha=0.7$. During the inference, we sampled $K\cdot N$ trajectories to accurately estimate the reaction probability. For USPTO-50k, we set $K=20$ while for, due to limited computational resources, we set $K=10$ for USPTO-MIT.
@@ Appendix B Generalization of Round-Trip Accuracy
We propose a generalization of round-trip accuracy that allows to use of a wider class of machine-learning models to assess the reaction feasibility. We call this metric Feasible Thresholded Count (FTC). For a single product, the top-k FTC value denotes the percentage of feasible reactions among top $k$ reactions returned by the model. The feasibility is estimated with an auxiliary model $\mathrm{RFM}$ described in further in this section. The exact formula for top-k FTC calculated on a product $p$ and retrosynthesis model $f$ is given by:
$\mathrm{F_{FTC}}(f,p,k)=\frac{1}{k}\sum_{i=1}^{k}\mathbb{1}[\mathrm{RFM}(f(p)_{i})\geq t],$ (4)
where $\mathrm{RFM}(f(p)_{i})\in[0,1]$ is the output of the reaction feasibility model for the $i$-th reaction proposed by $f$, and $t$ is a feasibility threshold given by the user. We assume that $\mathrm{RFM}(x)=1$ for reaction $x\in D$. We report the top-k FTC for the entire dataset $D$, which can be written as $FTC(f,k)=\frac{1}{n}\sum_{i}^{n}\mathrm{F_{FTC}}(f,p_{i},k)$.
@@ B.1 Reaction Feasibility Model (RFM)
Architecture. Our RFM implementation consists of two GINE [66] Graph Neural Networks (GNN) with a Transformer [52] layer and attention pooling at the top that creates product and reactant embeddings which are then concatenated and fed into the $\mathrm{MLP}$ layer.
Checkpoints for USPTO-50k. To train the model, we augmented the USPTO-50k dataset with negative (non-feasible) reactions using two methods: 1) application of existing forward templates to obtain a novel product from existing reactants, 2) swapping a product in the reaction with another product that is similar to the original one in terms of Tanimoto similarity. Such an approach ensured that the generated negative reactions are not trivially unfeasible (they use an existing template and/or the product is not strikingly different from the reactants), but still are very unlikely to occur in reality (the original reactants were reported to return a different product). We obtained a reaction feasibility dataset with a negative-to-positive ratio of 5:1. We trained two distinct checkpoints of feasibility models: RFM-Train-50k and RFM-Eval-50k. The RFM-Train was trained only on the train split of the reaction feasibility dataset and was then used to calculate the reward in the RetroGFN during the training.
@@ B.2 Experiments
We trained the RetroGFN using the RFM-Train model as a feasibility proxy and compared it on top-k accuracy and our FTC metric. We used the same hyperparameters as in Appendix A, but with $\mathrm{n\_dataset}=80,\mathrm{n\_replay}=16,\mathrm{n\_forward}=32$ and $\beta=12$. The results (Tables 5, 6, 7 and 8) mimic the ones from the main paper: our RetroGFN outperforms the model on FTC metric while obtaining competitive results on the standard top-k accuracy. The experiments show that our RetroGFN can leverage any machine-learning feasibility proxy. We believe that training a reliable and powerful feasibility proxy is a promising direction for future work.
@@
GLN 74.0 ± 43.9 57.4 ± 31.8 49.0 ± 28.2 38.1 ± 23.3 28.4 ± 18.9 16.4 ± 12.0
MEGAN 70.4 ± 45.7 54.2 ± 31.1 46.7 ± 27.4 37.3 ± 23.2 28.6 ± 18.9 18.5 ± 13.4
MHNreact 72.0 ± 44.9 54.1 ± 30.6 45.5 ± 26.3 35.0 ± 21.6 25.6 ± 17.2 14.3 ± 10.6
LocalRetro 73.4 ± 44.2 57.2 ± 30.6 49.8 ± 27.4 40.0 ± 23.6 31.4 ± 19.9 20.1 ± 13.6
RetroKNN 72.0 ± 44.9 54.7 ± 30.4 47.0 ± 26.6 35.7 ± 21.9 22.5 ± 15.6 9.5 ± 6.9
RootAligned 75.9 ± 42.8 57.2 ± 29.7 49.7 ± 27.2 40.3 ± 23.9 31.4 ± 19.9 16.1 ± 10.9
Chemformer 74.9 ± 43.4 41.5 ± 25.3 28.5 ± 19.5 15.7 ± 12.2 8.2 ± 7.2 3.4 ± 3.3
RetroGFN 72.4 ± 44.7 61.0 ± 31.9 54.1 ± 29.3 46.0 ± 26.6 38.9 ± 24.2 29.1 ± 19.5
@@
GLN 74.8 ± 43.4 64.2 ± 33.4 58.3 ± 30.2 49.7 ± 26.0 40.6 ± 21.9 27.6 ± 17.0
MEGAN 72.9 ± 44.5 61.5 ± 32.8 55.6 ± 29.3 47.2 ± 25.1 38.8 ± 21.2 27.1 ± 16.5
MHNreact 73.3 ± 44.2 61.7 ± 33.2 55.2 ± 29.5 46.1 ± 24.9 36.6 ± 20.9 22.7 ± 15.5
LocalRetro 76.3 ± 42.5 65.4 ± 32.2 59.4 ± 29.2 51.3 ± 25.4 42.9 ± 21.9 30.2 ± 16.9
RootAligned 77.6 ± 41.7 66.0 ± 31.5 60.0 ± 28.7 51.7 ± 25.3 42.4 ± 21.6 22.8 ± 12.7
Chemformer 73.8 ± 44.0 47.1 ± 29.5 34.2 ± 23.7 20.2 ± 16.3 11.0 ± 10.3 4.6 ± 5.0
RetroKNN 75.0 ± 43.3 62.9 ± 32.1 56.1 ± 28.7 45.1 ± 24.5 29.8 ± 18.9 12.8 ± 9.1
RetroGFN 77.4 ± 41.8 69.1 ± 32.5 64.5 ± 29.8 58.1 ± 26.8 51.7 ± 24.4 40.7 ± 20.2
@@ Appendix C Ablations
In this section we study a simple model-agnostic way of leveraging the Chemformer-Train to maximize the results of round-trip accuracy metric. The idea is to filter the results that are not backtranslated by the Chemformer-Train model during the evaluation. Tables 9 and 10 show that such an approach significantly improves the round-trip accuracy results, but with the costs of a drastic decrease of a standard top-k accuracy, especially for larger values of $k$.
@@
Table 10: Top-k round-trip accuracy results on USPTO-50k for models that use the Chemformer-Train filtering. The best results in every column and group are bolded. We observe that the performance of all the models is significantly improved, however with the price of degrading the top-k accuracy results.
@@ D.1 Round-trip Accuracy
The main limitation of the top-k round-trip accuracy is that it relies on the forward reaction prediction model which suffers both false negative and false positive errors. However, we believe that there is an inherent epistemic uncertainty within the notion of feasibility (we cannot screen all the reactions) and any sensible retrosynthesis metric will have some portion of false negatives (it will not take all feasible reactions into account). In comparison to top-k accuracy, our round-trip accuracy has a strictly lower number of false negatives, while keeping false positives on a decent level. We believe that the round-trip will benefit from the further improvements of the forward reaction prediction model and we leave it for future work.
@@ D.2 RetroGFN
Top-k Accuracy. The main limitation of our RetroGFN method is its results on top-k accuracy for $k<5$. At first glance, it looks like a trade-off necessary to achieve excellent results on the round-trip accuracy. We argue that it may be caused by two things: 1) other hyperparameters of the model are not optimal for top-k accuracy, 2) the GFlowNet framework struggles with spiky reward function, and 3) the parametrization of the composition process is sub-optimal. It is possible that further refinements of the method could improve the results.
Leveraging Chemformer-Train. The fact that RetroGFN leverages the Chemformer-Train checkpoint can be seen as an unfair advantage because a similar Chemformer-Eval model is used in the round-trip accuracy computation. However, we think that fairness comes from the fact that all models use the same data splits for training or evaluation. The models differ in the way they learn from the training data and leveraging the Chemformer-Train is yet another way of learning. It does not inject any new knowledge that cannot be extracted from the training data. Once the round-trip accuracy metric is established, it becomes reasonable to optimize it using Chemformer-Train. Moreover, we believe that Chemformer-Eval and Chemformer-Train are expected to be similar because they have similar goals: 1) to extract as much information from the train and test split as possible, and 2) to extract as much information from the train split as possible. It is sensible then that they share architecture. The difference should come from the data split used for training.
@@ Appendix E Computational Resources
We ran all the experiments on Nvidia V100 and A100 GPUs. The training of our model takes no more than 48h per checkpoint. When experimenting with the architecture and different feasibility proxy models, we trained no more than 100 checkpoints. For all the baselines, we used already trained checkpoints and only evaluated them on USPTO-50k and USPTO-MIT. The evaluation time depends on the model, but in total, it took no more than 400 GPU hours. It gives the upper bound of 5200 GPU hours for the total experimenting costs.

### v2

@@ 6 Conclusions
In this paper, we provided empirical arguments for the importance of reporting the round-trip accuracy in the single-step retrosynthesis model evaluation. Leveraging the GFlowNet framework, which is designed for tasks where plenty of sensible solutions are desired, we developed a RetroGFN model that achieves competitive results on top-k accuracy and on the top-k round-trip accuracy. We discuss the limitations of the paper in the D.
7 Future Work
There are two main directions of future work on RetroGFN. The first is to mitigate its current limitations including an improvement of the inference, adaptation of more robust feasibility models, or tuning the RetroGFN’s design choices. The second direction is to develop novel single-step retrosynthesis models that are guided by a feasibility model. An especially interesting idea is to design a template-free version of RetroGFN that would unlock the full potential of the GFlowNets framework to generate diverse reactions.
@@ Acknowledgements
The research of P. Gaiński was supported by the National Science Centre (Poland), grant no. 2022/45/B/ST6/01117. The research of M. Śmieja was supported by the National Science Centre (Poland), grant no. 2023/49/B/ST6
/01137. The research of J. Tabor was supported by the National Science Centre (Poland), grant no. 2021/41/B/ST6/01370. The research of M. Koziarski was supported by funding from CQDM Fonds d’Accélération des Collaborations en Santé (FACS) / Acuité Québec and Genentech. We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid (ACK Cyfronet Athena, HPC) for providing computer facilities and support within computational grant no PLG/2023/016550. For the purpose of Open Access, the author has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
Credit Statement
Piotr Gaiński: Conceptualization, Methodology, Investigation, Software, Writing - Original Draft. Michał Koziarski: Conceptualization, Methodology, Writing - Review & Editing. Krzysztof Maziarz: Conceptualization, Methodology, Writing - Review & Editing. Marwin Segler: Conceptualization, Methodology, Writing - Review & Editing. Jacek Tabor: Conceptualization, Writing - Review & Editing. Marek Śmieja: Supervision, Conceptualization, Writing - Review & Editing.
@@ Appendix A RetroGFN Details
All neural networks in RetroGFN used the same hidden dimension $h=200$. To obtain initial node and edge features for products, we used featurization from [kearnes2016molecular]. For the reactant pattern, we used the same edge featurization and a custom node featurization that accounted for atom type, degree, aromaticity, whether the atom was mapped in the original template, relative charge difference between product and reactant atom in the original template, and analogous implicit hydrogen difference. The node features for both products and reactant patterns were enriched with random walk positional encoding of size $\mathrm{n\_random\_walk}=16$.
Product node encoder $\mathrm{gnn}_{1}$ consists of $\mathrm{num\_layer\_1}=4$ layers of the MPNN convolution [gilmer2017neural] and one Transformer layer with $\mathrm{num\_heads}=8$. The reactant pattern encoder differs only in the number of layers $\mathrm{num\_layer\_2}=3$. Multi-layer perceptrons $\mathrm{MLP}_{1},\mathrm{MLP}_{2},\mathrm{MLP}_{3}$ had one hidden layer (with hidden dimension $h$) and used the GeLU activation function.
During training, we used a combination of three sampling methods: 1) standard exploratory sampling from the forward policy $P_{F}$ with some $\epsilon$ probability of taking random actions, 2) backward sampling from replay buffer [shen2023towards, fedus2020revisiting, jain2022biological], and 3) backward sampling from the dataset $D$. Backward sampling starts with a terminal state and samples the trajectory in the backward direction using the backward policy.
During the training probability of taking random action in the forward policy was set to $\epsilon=0.05$, the number of sampled forward trajectories in the batch was $\mathrm{n\_forward}=16$, and the analogous numbers for backward dataset trajectories and backward replay buffer trajectories were $\mathrm{n\_dataset}=96$ and $\mathrm{n\_replay}=16$. The model was trained with Adam optimizer [kingma2014adam] with a learning rate $lr=0.0005$ (with other parameters set to default values in the torch implementation) for $\mathrm{n_{i}terations}=25000$ iterations. In the evaluation, the forward policy temperature was set to $\alpha=0.7$. During the inference, we sampled $K\cdot N$ trajectories to accurately estimate the reaction probability. For USPTO-50k, we set $K=20$ while, due to limited computational resources, we set $K=10$ for USPTO-MIT.
@@ Appendix B Generalization of Round-Trip Accuracy
We propose a generalization of round-trip accuracy that allows to use of a wider class of machine-learning models to assess the reaction feasibility. We call this metric Feasible Thresholded Count (FTC). For a single product, the top-k FTC value denotes the percentage of feasible reactions among the top $k$ reactions returned by the model. The feasibility is estimated with an auxiliary model $\mathrm{RFM}$ described further in this section. The exact formula for top-k FTC calculated on a product $p$ and retrosynthesis model $g$ is given by:
$\mathrm{F_{FTC}}(g,p,k)=\frac{1}{k}\sum_{i=1}^{k}\mathbb{1}[\mathrm{RFM}(g(p)_{i})\geq t],$ (4)
where $\mathrm{RFM}(g(p)_{i})\in[0,1]$ is the output of the reaction feasibility model for the $i$-th reaction proposed by $g$, and $t$ is a feasibility threshold given by the user. We assume that $\mathrm{RFM}(x)=1$ for reaction $x\in D$. We report the top-k FTC for the entire dataset $D$, which can be written as $FTC(g,k)=\frac{1}{n}\sum_{i}^{n}\mathrm{F_{FTC}}(g,p_{i},k)$.
@@ B.1 Reaction Feasibility Model (RFM)
B.1.1 Architecture
Our RFM implementation consists of two Graph Neural Networks (GNN) layers with a Transformer [vaswani2017attention] layer and attention pooling at the top that creates product and reactant embeddings, which are then concatenated and fed into the $\mathrm{MLP}$ layer.
B.1.2 Checkpoints for USPTO-50k
To train the model, we augmented the USPTO-50k dataset with negative (non-feasible) reactions using two methods: 1) application of existing forward templates to obtain a novel product from existing reactants, 2) swapping a product in the reaction with another product that is similar to the original one in terms of Tanimoto similarity. Such an approach ensured that the generated negative reactions are not trivially unfeasible (they use an existing template and/or the product is not strikingly different from the reactants), but still are very unlikely to occur in reality (the original reactants were reported to return a different product). We obtained a reaction feasibility dataset with a negative-to-positive ratio of 5:1. We trained two distinct checkpoints of feasibility models: RFM-Train-50k and RFM-Eval-50k. The RFM-Train was trained only on the train split of the reaction feasibility dataset and was then used to calculate the reward in the RetroGFN during the training.
@@ B.2 Experiments
We trained the RetroGFN using the RFM-Train model as a feasibility proxy and compared it on top-k accuracy and our FTC metric. We used the same hyperparameters as in A, but with $\mathrm{n\_dataset}=80,\mathrm{n\_replay}=16,\mathrm{n\_forward}=32$ and $\beta=12$. The results (Tables 7, 8, 9 and 10) mimic the ones from the main paper: our RetroGFN outperforms the model on FTC metric while obtaining competitive results on the standard top-k accuracy. The experiments show that our RetroGFN can leverage any machine-learning feasibility proxy. We believe that training a reliable and powerful feasibility proxy is a promising direction for future work.
@@
GLN 74.0 57.4 49.0 38.1 28.4 16.4
MEGAN 70.4 54.2 46.7 37.3 28.6 18.5
MHNreact 72.0 54.1 45.5 35.0 25.6 14.3
LocalRetro 73.4 57.2 49.8 40.0 31.4 20.1
RetroKNN 72.0 54.7 47.0 35.7 22.5 9.5
RootAligned 75.9 57.2 49.7 40.3 31.4 16.1
Chemformer 74.9 41.5 28.5 15.7 8.2 3.4
RetroGFN 72.4 61.0 54.1 46.0 38.9 29.1
@@
GLN 74.8 64.2 58.3 49.7 40.6 27.6
MEGAN 72.9 61.5 55.6 47.2 38.8 27.1
MHNreact 73.3 61.7 55.2 46.1 36.6 22.7
LocalRetro 76.3 65.4 59.4 51.3 42.9 30.2
RootAligned 77.6 66.0 60.0 51.7 42.4 22.8
Chemformer 73.8 47.1 34.2 20.2 11.0 4.6
RetroKNN 75.0 62.9 56.1 45.1 29.8 12.8
RetroGFN 77.4 69.1 64.5 58.1 51.7 40.7
@@ Appendix C Ablations
In this section, we study a simple model-agnostic way of leveraging the Chemformer-Train to maximize the results of the round-trip accuracy metric. The idea is to filter the results that are not backtranslated by the Chemformer-Train model during the evaluation. Tables 11 and 12 show that such an approach significantly improves the round-trip accuracy results, but with the cost of a drastic decrease in a standard top-k accuracy, especially for larger values of $k$.
@@
Table 12: Top-k round-trip accuracy results on USPTO-50k for models that use the Chemformer-Train filtering. The best results in every column and group are bolded. We observe that the performance of all the models is significantly improved, however, at the price of degrading the top-k accuracy results.
@@ D.1 Round-trip Accuracy
The main limitation of the top-k round-trip accuracy is that it relies on the forward reaction prediction model, which suffers from both false negative and false positive errors. However, we believe that there is an inherent epistemic uncertainty within the notion of feasibility (we cannot screen all the reactions), and any sensible retrosynthesis metric will have some portion of false negatives (it will not take all feasible reactions into account). In comparison to top-k accuracy, our round-trip accuracy has a strictly lower number of false negatives, while keeping false positives at a decent level. We believe that the round-trip will benefit from the further improvements of the forward reaction prediction model and we leave it for future work.
@@ D.2 RetroGFN
D.2.1 Top-k Accuracy
The main limitation of our RetroGFN method is its results on top-k accuracy for $k<5$. At first glance, it looks like a trade-off necessary to achieve excellent results on the round-trip accuracy. We argue that it may be caused by two things: 1) other hyperparameters of the model are not optimal for top-k accuracy, 2) the GFlowNet framework struggles with a spiky reward function, and 3) the parametrization of the composition process is sub-optimal. It is possible that further refinements of the method could improve the results.
D.2.2 Leveraging Chemformer-Train
The fact that RetroGFN leverages the Chemformer-Train checkpoint can be seen as an unfair advantage because a similar Chemformer-Eval model is used in the round-trip accuracy computation. However, we think that fairness comes from the fact that all models use the same data splits for training or evaluation. The models differ in the way they learn from the training data, and leveraging the Chemformer-Train is yet another way of learning. It does not inject any new knowledge that cannot be extracted from the training data. Once the round-trip accuracy metric is established, it becomes reasonable to optimize it using Chemformer-Train. Moreover, we believe that Chemformer-Eval and Chemformer-Train are expected to be similar because they have similar goals: 1) to extract as much information from the train and test split as possible, and 2) to extract as much information from the train split as possible. It is sensible then that they share architecture. The difference should come from the data split used for training.
@@ Appendix E Computational Resources
We ran all the experiments on Nvidia V100 and A100 GPUs. The training of our model takes no more than 48h per checkpoint. When experimenting with the architecture and different feasibility proxy models, we trained no more than 100 checkpoints. For all the baselines, we used already trained checkpoints and only evaluated them on USPTO-50k and USPTO-MIT. The evaluation time depends on the model, but in total, it took no more than 400 GPU hours. It gives the upper bound of 5200 GPU hours for the total experiments’ costs.